# Semi-Supervised Deep Sobolev Regression: Estimation, Variable Selection and Beyond

### v1

@@ Appendix E Proofs in Results in Section 6
Proof of Corollary 6.4.
By a same argument as the proof of Theorem 5.4, we have
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\nabla(\widehat{u}_{{\mathcal{D}}}^{\lambda}-u_{*})\|_{L^{2}(\mu_{X})}\Big{]}\leq{\mathcal{O}}\Big{(}\Big{(}1+\sigma^{-\frac{3s}{2(d+4s)}}\Big{)}n^{-\frac{s}{2(d+4s)}}\Big{)}.$ (65)
Since that $(u_{*},f_{*})$ satisfies (13), by Evans (2010, Section 6.1), the following variational equation holds
$(f_{*},v)_{L^{2}(\mu_{X})}=(\nabla u_{*},\nabla v)_{L^{2}(\mu_{X})}+(wu_{*},v)_{L^{2}(\mu_{X})},\quad\forall v\in H^{1}(\mu_{X}).$ (66)
Subtracting (66) from (16) deduces the following equality for each $v\in H^{1}(\mu_{X})$
$(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{*},v)_{L^{2}(\mu_{X})}=(\nabla(\widehat{u}_{{\mathcal{D}}}^{\lambda}-u_{*}),\nabla v)_{L^{2}(\mu_{X})}+(w(\widehat{u}_{{\mathcal{D}}}^{\lambda}-u_{*}),v)_{L^{2}(\mu_{X})}.$
According to Cauchy-Schwarz inequality, it follows that for each $v\in H^{1}(\mu_{X})$
$|(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{*},v)_{L^{2}(\mu_{X})}|$
$\leq\|\nabla(\widehat{u}_{{\mathcal{D}}}^{\lambda}-u_{*}\|_{L^{2}(\mu_{X})}\|\nabla v\|_{L^{2}(\mu_{X})}+\|w\|_{L^{\infty}(\mu_{X})}\|\widehat{u}_{{\mathcal{D}}}^{\lambda}-u_{*}\|_{L^{2}(\mu_{X})}\|v\|_{L^{2}(\mu_{X})}$
$\leq C\Big{\{}\|\nabla(\widehat{u}_{{\mathcal{D}}}^{\lambda}-u_{*})\|_{L^{2}(\mu_{X})}+\|\widehat{u}_{{\mathcal{D}}}^{\lambda}-u_{*}\|_{L^{2}(\mu_{X})}\Big{\}}\|v\|_{H^{1}(\mu_{X})},$
which implies
$\|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{*}\|_{(H^{1}(\mu_{X}))^{*}}\leq C\Big{\{}\|\nabla(\widehat{u}_{{\mathcal{D}}}^{\lambda}-u_{*})\|_{L^{2}(\mu_{X})}+\|\widehat{u}_{{\mathcal{D}}}^{\lambda}-u_{*}\|_{L^{2}(\mu_{X})}\Big{\}}.$
Taking expectation with respect to ${\mathcal{D}}\sim\mu^{n}$ yields
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{*}\|_{(H^{1}(\mu_{X}))^{*}}\Big{]}\leq C{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\|\nabla(\widehat{u}_{{\mathcal{D}}}^{\lambda}-u_{*})\|_{H^{1}(\mu_{X})}\Big{]}.$ (67)
Combining (65) and (67) completes the proof. ∎
@@ Appendix F Additional Experiments Results
F.1 Additional Examples for Derivative Estimation
Example F.1.
@@ F.1 Additional Examples for Derivative Estimation
We consider a toy problem in two-dimensions, where the support of the marginal distribution $\mu_{X}$ approximately coincides with the coordinate subspace $[0,1]\times\{0\}$. Precisely the first element of the covariate is uniformly sampled from $[-1,1]$, whereas the second one is drawn from a normal distribution $N(0,0.05)$. The underlying regression function is $f_{0}(x)=x_{1}^{2}$, and labels are generated by $Y=f_{0}(X)+\xi$, where the noise term $\xi\sim N(0,0.1)$. The regularization parameter is set as $\lambda=1.0\times 10^{-4}$.
@@
In all cases the accuracy on the $\operatorname{supp}(\mu_{X})$ is high, see Figure 3 (top), but the least square regressor fails to extend the approximation and smoothness outside the support, as the least square loss is insensible to errors out of $\operatorname{supp}(\mu_{X})$. While the landscape of DORE is smoother compared to the least square regressor, SDORE further extends the smoothness to $[-1,1]^{2}$ as it utilizes unlabeled samples from $\nu_{X}$.
@@
We examine the partial derivative estimation with respect to $x_{1}$ and $x_{2}$ on $[-1,1]^{2}$ and display the result in Figure 3. As expected, the least square one is unstable compared to the DORE and SDORE. Also, we can tell from the bottom right of Figure 3 that $x_{2}$ is the irrelevant variable.
@@
Figure 3: Effect of the regularization technique on function fitting for a toy problem $f_{0}(x)=x_{1}^{2}$. (a) Landscape of the primitive function and its partial derivatives. The train samples are plotted in black dots. (b) Least square fitting estimation. (c) DORE estimation. (d) SDORE estimation.
F.2 Additional Examples for Variable Selection
Example F.2 (Simulation Studies).
@@ F.2 Additional Examples for Variable Selection
Consider $f_{0}(x)=2x_{1}^{2}+e^{x_{2}}+2\sin(x_{3})+2\cos(x_{4}+1)$, with observations $y=f_{0}(x)+\xi$, where $x\in\mathbb{R}^{10}$ and $\xi$ is a white noise, sampled from a normal distribution with the signal to noise ratio to be 25. $x_{1}$ to $x_{4}$ are drawn from the uniform distribution on $[0,1]$, and the rest noise variables are drawn from the uniform distribution on $[0,0.05]$. The regularization parameter $\lambda$ is set as $1.0\times 10^{-4}$ for DORE and SDORE.
@@
We repeat the process for least square regressor (LS), DORE and SDORE respectively, and evaluate the models on 100 different test sets with sample size 500. We report the partial derivative of three estimators with respect to $x_{i}$, the selection error (mean of false positive rate and false negative rate) the prediction errors (RMSE) in Figure 4. The results indicate least square fails to identify the correct dependent variables. DORE and SDORE yield smaller prediction error, while pointing out that $x_{1}$ to $x_{4}$ are the relevant variables.
@@
Figure 4: (left) Partial derivatives estimation on a variable selection problem in $\mathbb{R}^{10}$ where $f_{0}$ is dependent on the $x_{1}$ to $x_{4}$. The dashed line is the threshold for variable selection. (center) Selection error on test set. (right) Prediction error on test set.
Example F.3 (California Housing Dataset).
We explore the regression task on California housing dataset. The dataset consists of 8 feature variables: MedInc (median income in block group), HouseAge (median house age in block group), AveRooms (average number of rooms per household), AveBedrms (average number of bedrooms per household), Population (block group population), AveOccup (average number of household members), Latitude (block group latitude), Longitude (block group longitude); and the target variable: MedHouseVal (median house value).
We manually append 7 irrelevant noise variables $\{\text{noise}_{1}\}_{i=1}^{n}$ from the uniform distribution on $[0,1]$ to the features, and then carry out the Sobolev penalized process for $\lambda=0.001,0.01$ and 0.1, respectively. We report the partial derivatives with respect to feature variables and prediction error in Figure 5. All experiments succeed ruling out appended noise features and indicate that AveOccup, Latitude, Longitude and MedInc are the most dominant variables.
Figure 5: DORE on California housing dataset. (left) Partial derivatives with respect to feature variables in descending order. (right)Prediction error on test set.
F.3 Additional Results for Example 6.3
Experiments setup We approximate the unknown solution using a neural network, which has two hidden layers, and each layer containing 64 units. In this experiment, we use the Adam optimizer (Kingma and Ba, 2015) with a learning rate $1.0\times 10^{-3}$. We use 10000 sample points and set the batch size as 128. We set the number of epochs as 1000.
In the case where $\sigma=20\%$ and $\lambda=1.0\times 10^{-4}$, we give a more intuitive comparison between the ground truth solution and their reconstructions, we also show the point-wised absolute error in Figure 6. This figure illustrates the high accuracy of the reconstructions even in the presence of high measurement noise. In Figure 7, we show a comparison between the true source term and the inverted source term, illustrating that our inverted source term algorithm does not degrade significantly in accuracy as the noise gets larger.
In the case where $\sigma=20\%$ and $\lambda=1.0\times 10^{-4}$, we provide a more intuitive comparison between the ground truth solution and its reconstruction, along with the point-wise absolute error, as shown in Figure 6. This figure clearly demonstrates the high accuracy of the reconstructions, even in the presence of high measurement noise. Additionally, in Figure 7, we present a comparison between the true source term and the recovered source term, illustrating that our SDORE-based approach for recovering the source term maintains a high level of accuracy even as the noise level increases.
Refer to captionRefer to captionRefer to caption
(a) Solution $u$
Refer to captionRefer to captionRefer to caption
(b) Derivative $D_{1}u$
Refer to captionRefer to captionRefer to caption
(c) Derivative $D_{2}u$
Figure 6: The ground truth $(f_{*},u_{*},D_{1}u_{*},D_{2}u_{*})$ (top), the reconstructions $(\widehat{f}_{{\mathcal{D}}}^{\lambda},\widehat{u}_{{\mathcal{D}}}^{\lambda},D_{1}\widehat{u}_{{\mathcal{D}}}^{\lambda},D_{2}\widehat{u}_{{\mathcal{D}}}^{\lambda})$ (middle), and the point-wise absolute error of the reconstructions (bottom) under the noise level $\sigma=20\%$ and regularization parameter $\lambda=1.0\times 10^{-4}$ in Example 6.3.
Refer to captionRefer to caption
(a) Clear data
Refer to captionRefer to caption
(b) Noisy data
Refer to captionRefer to caption
(c) Exact source
Refer to captionRefer to caption
(d) Recovery
Refer to captionRefer to caption
(e) Point-wise error
Figure 7: The ground truth solution $u_{*}$ (clear data), the measurement $Y$ (noisy data), the exact source $f_{*}$, the recovery of source $\widehat{f}_{{\mathcal{D}}}^{\lambda}$ and the piece-wise absolute value error $|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{*}|/|f_{*}|$ in Example 6.3. $\delta=10\%$ and $\lambda=1.0\times 10^{-8}$ (top) $\delta=20\%$ and $\lambda=1.0\times 10^{-4}$ (bottom)

### v2

@@ Appendix F Additional Experiments Results
F-A Additional Examples for Derivative Estimation
Example 3.
@@ F-A Additional Examples for Derivative Estimation
We consider a toy problem in two-dimensions, where the support of the marginal distribution $\mu_{X}$ approximately coincides with the coordinate subspace $[0,1]\times\{0\}$. Precisely the first element of the covariate is uniformly sampled from $[-1,1]$, whereas the second one is drawn from a Gaussian distribution $N(0,0.05)$. The underlying regression function is $f_{0}(x)=x_{1}^{2}$, and labels are generated by $Y=f_{0}(X)+\xi$, where the noise term $\xi\sim N(0,0.1)$. The regularization parameter is set as $\lambda=1.0\times 10^{-4}$.
@@
In all cases the accuracy on the $\operatorname{supp}(\mu_{X})$ is high, see Figure 3 (top), but the least-squares regressor fails to extend the approximation and smoothness outside the support, as the least-squares loss is insensible to errors out of $\operatorname{supp}(\mu_{X})$. While the landscape of DORE is smoother compared to the least-squares regressor, SDORE further extends the smoothness to $[-1,1]^{2}$ as it utilizes unlabeled samples from $\nu_{X}$.
@@
We examine the partial derivative estimation with respect to $x_{1}$ and $x_{2}$ on $[-1,1]^{2}$ and display the result in Figure 3. As expected, the least-squares one is unstable compared to the DORE and SDORE. Also, we can tell from the bottom right of Figure 3 that $x_{2}$ is the irrelevant variable.
@@
Figure 3: Effect of the regularization technique on function fitting for a toy problem $f_{0}(x)=x_{1}^{2}$. (a) Landscape of the primitive function and its partial derivatives. The train samples are plotted in black dots. (b) least-squares fitting estimation. (c) DORE estimation. (d) SDORE estimation.
F-B Additional Examples for Variable Selection
Example 4.
@@ F-B Additional Examples for Variable Selection
Consider the regression function $f_{0}(x)=2x_{1}^{2}+e^{x_{2}}+2\sin(x_{3})+2\cos(x_{4}+1)$, with observations $Y=f_{0}(X)+\xi$, where $X\in\mathbb{R}^{10}$ and $\xi$ is a white noise, sampled from a Gaussian distribution with the signal to noise ratio to be 25. The first four elements of $X$ are drawn from the uniform distribution on $[0,1]$, and the rest noise variables are drawn from the uniform distribution on $[0,0.05]$. The regularization parameter $\lambda$ is set as $1.0\times 10^{-4}$ for SDORE.
@@
We repeat the process for least-squares regressor and SDORE, respectively, and evaluate the models on a test set with sample size 1000. We report the estimated mean square of partial derivative by both estimators with respect to $x_{i}$, the mean selection error (mean of false positive rate and false negative rate) the root mean squared prediction error on the primitive function in Figure 4. The results indicate least-squares regression fails to identify the correct dependent variables and has larger prediction error. In contrast, SDORE yields smaller prediction error, and points out that $x_{1}$ to $x_{4}$ are the relevant variables.
@@
Figure 4: (left) Empirical mean square of the partial derivatives estimated by least-squares regression (LS) and SDORE on a variable selection problem in $\mathbb{R}^{10}$ where $f_{0}$ is dependent on the $x_{1}$ to $x_{4}$. The dashed line is the 75 % quantile threshold for variable selection. (center) Mean variable selection error for the estimated derivative function on test set. (right) Root mean squared prediction error for the primitive function on test set.
Acknowledgment
The authors would like to express their sincere gratitude to the Editor-in-Chief, Associate Editor, and all reviewers for their careful review of the manuscript. Their insightful comments and constructive suggestions have significantly improved the quality and clarity of this paper. This work is supported by the National Key Research and Development Program of China (No. 2024YFA1014202), by the National Natural Science Foundation of China (No. 12125103, No. U24A2002, No. 12371441), and by the Fundamental Research Funds for the Central Universities. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.